« November 2005 | Main | January 2006 »

December 31, 2005

Project Status: 31 December 2005 (22 UTC)

We are now well into the post-frontal stage of this wet event. Precipitation is still falling in the mountains in a showery mode. Accumulations continue to increase; Blue Canyon has had about 10 inches since 05 UTC 30 December. There are lots of flooding issues throughout the region. Flooding in the lee of the Sierra is perhaps the most severe, with many parts of Reno under water. Also, a debris flow near Floriston (close to the Nevada-California state line) has closed down Interstate-80, perhaps for 2 or more days.

The SMART-R ceased operations at about 20 UTC so that the operator (Ken Howard) could get some rest in advance of the Sunday-Monday precipitation event. There are some hardware issues on the SMART-R data acquistion computer that will need to be addressed before data collection can begin again. This will be addressed first thing Sunday morning.

XPOL ceased operations at about 21 UTC so that the engineer (Kurt Clark) could address some weatherproofing issues on parts of the radar exposed to the elements.

The last supplemental sounding at Oakland was at 21 UTC

The last sounding at the Sloughhouse field site was at 18 UTC, just after the cold front passed the site. Many of the soundings at Sloughhouse experienced a porpoising effect, where ice builds up on the balloon and inhibits ascent. These sondes only rose to about 3 km before descending and slamming into terrain. Attempts will be made to overcome this problem.

IOP 4 will be declared over at 00 UTC 1 January. We anticipate that IOP 5 will begin at 00 UTC 2 January and last about 24 hours. Supplemental soundings at Oakland have been requested at 03, 06, 09, 15, 18 and 21 UTC 2 January. The current plan is for Sloughhouse sounding operations to begin at 00 UTC 2 January (i.e., 3 pm PST 1 January release) and continue through 00 UTC 3 January (i.e., 3 pm PST 2 January release). Subtle adjustments to these plans may be made mid-day Sunday after further guidance is available.


Kingsmill

Project Status: 31 December 2005 (05 UTC)

Precipitation continues to fall steadily in the ARB. Over the last 8 hours, Blue Canyon and Alta have received about 2.5 inches of rain. XPOL continues to operate without any apparent problems. The 00 UTC sounding at Sloughhouse was partially successful; data were collected up to almost 500 mb before GPS lock was lost. Soundings will continue through the night and into the daytime hours Saturday. Supplemental soundings from Oakland will also continue.

Kingsmill

December 30, 2005

Project Status: 30 December 2005

After a relatively slow start this morning, precipitation in our target area has increased in intensity and areal coverage over the last few hours (since about 19 UTC). Precip rates have increased substantially at Blue Canyon and Alta since 19 UTC. IOP 4 began this morning at 12 UTC and is likely to continue into tomorrow afternoon. We will discuss the specific IOP end time on Saturday during the daily coordination call, but my guess is that it will be around 00 UTC 1 January (4 pm PST Saturday). Here are some status updates:

XPOL: The replacement frequency-converter arrived this morning and was installed. After fixing a few other unrelated problems, the radar came back up and has been operating since about 21 UTC. The radar will be monitored closely this afternoon and tonight, with the hope that no further problems materialize. The ESRL/PSD engineering staff, particularly Kurt Clark and Tom Ayers, did a great job in troubleshooting the problem, rapidly shipping the replacement parts and executing the repair. Kurt and Tom, thanks for your efforts.

SMART-R: After some initial issues with the truck being stuck in the mud, Ken Howard was able to begin scanning operations at about 18 UTC.

Oakland Supplemental Sondes: The first supplemental launch was at 15 UTC this morning. The plan is for these 3-hourly supplemental launches to continue through 21 UTC Saturday.

Sloughhouse Sondes: Attempts were made to launch sondes at the Sloughhouse field site at 18 UTC and 21 UTC, but equipment problems precluded their execution. However, Dave White tells me that these problems have now been resolved and there should be a successful 00 UTC sonde (3 pm PST release). The plan is for these 3-hourly launches to continue into Saturday mid-day (subject to how long Dave White can perform the sounding operations by himself).


Kingsmill

December 29, 2005

Project Status: 29 December 2005

An extremely wet event is approaching our target area. Precipitation is likely to begin in earnest between 12-15 UTC Friday.

SMART-R operations will begin around16 UTC and will involve only 1 operator (Ken Howard). He will attempt to operate the radar by himself thru Friday and into Saturday morning.

XPOL is still down and will not be back up until at least Friday afternoon. The lead engineer for XPOL discovered that the frequency converter (or motor generator) for the radar failed and needs to be replaced. A replacement is being shipped to Auburn overnight and should arrive at the site around 11 am PST Friday. Installation of the replacement frequency converter may be difficult due to the expected inclement weather conditions. An update to XPOL status may be available during our Friday coordination call, but we probably won't know much until later Friday afternoon.

The Oakland rawinsonde equipment has been repaired and we have requested supplemental soundings at 3-hourly intervals to begin at 09 UTC Friday. We may not have provided enough lead time to start the supplementals at 09 UTC, so it is possible the first supplement launch will be 15 UTC Friday. We have asked for the supplemental launches to continue through 18 UTC Saturday. This end time might be adjusted during our Friday coordination call based on updated information.

We still have plans to launch sondes Friday at the Sloughhouse field site. We are hoping for a first launch at 9 am PST, but logistics issues may delay the first launch until 12 noon PST. The desire is to launch at 3-hourly intervals for as long as Dave White and the student trainees are capable, hopefully into late Friday night. An update on the Sloughhouse sounding status should be available during the Friday coordination call (Dave, give me a call in the morning when you get a chance). These launch times are meant to coincide with the Oakland rawinsonde launches. Note: a 21 UTC (1 pm PST) sonde at Oakland is actually released about an hour before (i.e, about 12 noon PST)

If the Oakland supplementals are able to start at 09 UTC, then we will declare that the start time for IOP 4. If not, then we will declare 12 UTC the start time for IOP 4. We anticipate that this IOP will last into mid-day Saturday.

Kingsmill

December 28, 2005

Project Status: 28 December 2005

The forecast from yesterday remains on track, with a slight delay in the expected start of precipitation in the American Basin. At present, it appears that only light precipitation will affect the ARB during late Thursday night into early Friday morning. Precipitation intensity should increase after daybreak on Friday. The XPOL crew (Sergey Matrosov) is already on site in Auburn ready to operate the radar for the event. I have been unable to communicate with NSSL today, but my understanding from talking with Dave Jorgensen yesterday was that at least one NSSL crew memeber would deploy to the ARB for operation of the SMART-R. Specific timing for initiation of scanning radar operations will be discussed during the coordination call Thursday.

XPOL status:

Upon arrival at Auburn today, Sergey Matrosov encountered problems with bringing up XPOL. After 1-2 hours of telephone consultation with Kurt Clark, the lead engineer for XPOL, the problems were not able to be solved. Therefore, Kurt will be traveling to Auburn on Thursday in an attempt to investigate and solve the problems. An update on XPOL status will not be available until Thursday mid-afternoon (i.e., after the mid-day coordination call).

Kingsmill

December 27, 2005

Project Status: 27 December 2005

HMT-West 2006 has resumed operation after our holiday break. We have missed a few events, but the forecast over the next week suggests that the wet weather will continue for a while. After the event occuring today into early tomorrow, the next event is forecast to start affecting the American River Basin late Thursday or early Friday. Therefore, crews from ESRL and NSSL will be deploying to the American River Basin to be ready for scanning radar data collection by late Thursday.

Sounding Status:

The upcoming event (IOP 4) will be our first opportunity to launch sondes at the Sloughhouse field site. Due to some setup and training issues, we will not be able to start sounding operations until Friday morning. Also, continuous (i.e.,24 h) sounding operations will not be possible for this IOP (however, we anticipate that continuous, 24 h sounding operations will be possible in subsequent IOP's). For IOP 4, we would like 3-hourly releases starting as early as possible on Friday (I have been told that this time is most likely to be 12 noon PST, but might be 9 am PST if a lot of other things go smoothly). These soundings would continue on a 3-hourly interval for as long as Dave White and the student trainees can safely operate without a break, which is likely to be into Friday evening local time.

We would like to ask for supplemental NWS soundings from OAK for this event. However, the sounding equipment at OAK is under repair at the present time. The timeline for repair of this equipment is not yet known. An update will be available tomorrow.

Instrument Status:

Surface meteorology instrumentation at Chico (CCO), Colfax (CFC) and Foresthill (FHL) is not updating on the HMT Network page. There is a very high probability that these problems are related to data communication issues from the site back to Boulder. In other words, the data is most likely being recorded properly at the site, but is not being telemetered back to Boulder. These problems will be addressed by ESRL staff over the next few days.

Kingsmill

December 22, 2005

Project Status: 22 December 2005 (Thursday morning)

IOP-3 has ended by default. Although more rain is on the way, the research radar crews must depart to catch their flights home for Christmas. The X-band stopped recording at 10:11 UTC, 22DEC; SMART-R continued to 12 UTC. Blue Canyon had 3.6 inches of rain in the 24 hours ending at 12 UTC and Auburn had 1.3 inches. Another good weather event - this one with lots of convective storms.

Next conference call Dec. 27th.

- Martner

Project Status: 21 December 2005 (Wednesday evening)

IOP-3 continues with both research radars scanning. This afternoon's convective storms with 55 dBZ cores in the foothills made the Oklahoma crew feel at home. Storm-total rain accumulations exceed 2 inches at Blue Canyon and 1.2 inches at Auburn. The Central Valley has had much less. NEXRAD shows one final wide swath of rain which should pass the foothills by midnight.

- Martner

December 21, 2005

Project Status: 21 December 2005 (Wednesday noon)

IOP-3 continues with SMART-R and X-Pol research radars still scanning away. A mid-day lull in the rain is expected to end by about 4 pm this afternoon, as the next big wave of moisture reaches the ARB. Water vapor is 250% of normal on this morning's sounding at Oakland. NEXRAD shows the approaching wave of rainy weather is already onshore as far south as Sonoma County. Forecasters are expecting 2 inches or more of new rain in the ARB from this afternoon to sunrise, and still more duing the day Thursday. Heaviest rain period should be from about 7 pm tonight to 1 am (PST) tomorrow morning. Longer range holds potential for wet weather as, or soon after, the project resumes, following the holiday break. See "Forecast" for details.

Radar crews will operate all day and evening. Regardless of the weather conditions, the research radars will be shut down no later than 4 a.m. Thursday, so that crews can catch their flights home to Oklahoma and Colorado for Christmas.

Next conference call: usual time on Tuesday, Dec. 27th.

- Martner

Project Status: 21 December 2005 (Wednesday morning)

After Tuesday's slow start, IOP-3 has shifted into a higher gear as thunderstorms and rainbands move into the foothills. Rain totals since yesterday afternoon include 3/4 inch at Auburn and 0.9 inches at Blue Canyon. Both research radars operated through the night and will continue all day and into tonight. Satellite IR images indicate much more rain will be coming to the ARB during the daylight period later today. The SSMI water vapor image from 13 UTC shows two atmospheric rivers of vapor converging on the California coast; one passes north of Hawaii and one is from south of Hawaii. S-band profiler data along the slope on I-80 occassionally show the melting layer radar brightband at about 3 km MSL, but often no brightband is aparent, perhaps because of the convection, or because of a warm-rain-like process.

Displays of HMT profilers in/near the HMT are now online on the HMT06 project web page. These include the 915-MHz wind profiler at Slough House (SHS) and the S-band precip profilers at Alta (ATA) and Colfax (CFX). Click on "HMT Network" on the page's left sidebar, then click on the appropriate dot on the Basin-Scale map. The realtime comparison of the low sweeps from the X-band radar and KDAX is also on line under "X-pol andKDAX Radar Intercomparison"

- Martner

Project Status: 20 December 2005 (Tuesday night)

IOP-3 is off to a very slow start. Only about 0.1 inches of rain fell in the ARB today and this evening - mostly from scattered showers at the southern tip of the precip swath, which now exends northwestward to Idaho. However, the IR imagery shows the next wave of high cloud has reached the coastline at Eureka. This oncoming system extends much farther south and seems unlikely to miss the ARB. We expect a lull of a few hours berfore it begins to affect the ARB. SMART-R will contine to scan and document the remants of the Tuesday wave and the onset of the Wednesday wave. The X-band will also continue overnight operations (in unmanned mode). The Tuesday-Thursday action will all be considered part of IOP-3.

- Martner

December 20, 2005

Project Status: 20 December 2005 (Tuesday noon)

IOP-3 is underway with both research radars scanning. Regional NEXRADs show precipitation moving across California's northern counties, but it appears to be diminishing and might well miss the ARB. Hence, expectations for wet weather today in the ARB have plummeted. Nevertheless, SMART-R and X-Pol will monitor at least into early evening, at which time, they might take a break, based on conditions then. Tomorrow looks much better as a more vigorous and extensive storm moves on shore (see "Forecast" for details). Radar crews should be prepared to operate all day Wednesday and late into Wednesday night.

Tomorrow's conference call will be the last one until Dec.27th.

- Martner

Project Status: 20 December 2005 (Tuesday morning)

Rain from the approaching storm has moved onshore as far south as Sonoma County by sunrise. IOP-3 will begin at 18 UTC today (10 a.m PST) when both research radars will begin scanning. Sacramento and Beale NEXRAD should be set to VCP-12.

- Martner

December 19, 2005

Project Status: 19 December 2005 (Monday noon)

IOP-2 ended this morning. It was a pretty respectable event with a total accumulation of about 4 inches of liquid equivalent at Blue Canyon, which was much more than originally predicted.

The next storm (a warm one) is forecast to move into the ARB around midday tomorrow. See "Forecast" for details. The research radar crews should be ready to start scanning by mid-morning on Tuesday and be prepared to keep going into Wednesday night.

Crews will return to Oklahoma and Colorado on Thursday, as originally scheduled.

- Martner

Project Status: 19 December 2005 (Monday mid-morning)

IOP-2 ends. We are taking advantage of the lengthy dry period this morning to declare an end to IOP-2. Both research radars have stopped scanning at 17:20 UTC. Possibilities for starting an IOP-3 in the near future will be discussed on today's conference call.

- Martner

Project Satus: 19 December 2005 (Monday morning)

The weekend storm has no clear back edge as a series of waves parade through in rapid succession. After an evening lull, another half inch of rain fell in the foothills last night and Blue Canyon picked up another inch. (Total at Blue Canyon since Saturday night is now 4.4 inches liq. equiv.) Heaviest action is now farther north however.

Research radars will continue operating at least until mid day.

- Martner

Project Status: 19 December 2005 (Sunday night)

IOP-2 is still in progress. Storm-total precipitation has reached about 2 inches in the Central Valley and foothills and more than 3 inches at Blue Canyon. However, precip has been much lighter and spotty this evening. Satellite imagry shows another wave of cold cloud approaching the shoreline near Point Arena. We will continue operating the research radars to monitor the lingering light precip of the Sunday storm and (we hope) to catch the anticipated intensification with the arrival of the next wave before sunrise Monday.

- Martner

December 18, 2005

Project Status: 18 December 2005 (Sunday noon)

This weekend storm has already produced more precip than was expected 2 days ago, with 2 inches of liq. equiv. at Blue Canyon and about 1.3 inches in the foothills so far. And the short-term forecasts are calling for another inch during the afternoon and evening, making this a respectable IOP event. Thereafter, a series of little impulses is predicted for Monday, Tuesday, Wednesday, but these appear to take aim somewhat north of the ARB. See "Forecast" for details.

The research radars will continue scanning; crews should expect to keep going at least to sunrise Monday.

Today's conference call did not get activated by the leader. We need to get that back on track tomorrow. And, on tomorrow's call, we will consider if the early to mid-week impulses are worthy of keeping the crews out here.

- Martner

Project Status: 18 December 2005 (Sunday, sunrise)

IOP-2 continues with good rainfall intensities in foothills. About 0.75-1.0 inches of liquid equivalent fell overnight in the ARB (rain in the foothills, snow at Blue Canyon). A melting layer brightband was prominent at about 2.4 km above Auburn earlier when the echoes were deep. More recently, the echoes have broken into two or more vertically separated layers that appear to merge higher up the barrier. Warm front appears to have moved into the foothills around 13 UTC, based on pressure trends. Surface winds are still from SSE at Auburn and gusting to 20 kts. VAD scans at Auburn and profilers in Central Valley show winds veer rapidy with height and come from west of south above about 1 km.

- Martner

Project Status: 17 December 2005 (Saturday evening)

IOP-2 is underway as both research radars monitor the light precipitation of the early stages of the oncoming storm. The X-band radar began recording data at 00:56 UTC (18DEC) from Auburn, and SMART-R began at 03:26 from Foresthill.

- Martner

December 17, 2005

Project Status: 17 December 2005 (Saturday noon)

The forecast models have taken a more encouraging tone for this weekend's storm, with the heavier precip beginning in the ARB around midnight PST tonight and continuing at good intensity through the daylight period on Sunday, then tapering off and ending by early Monday morning. Storm-total (liquid) ccumulations of 2 inches now seem likely. We will begin operating the research radars by 6 pm PST today to catch the early (weaker) stages of the storm and follow its evolution. Research radar crews should expect about 36 hours of continuous operations at this point. There will be about a 1 day break after this storm before the next one arrives, altough that one is looking weaker in the models than it did yesterday. See "Forecast" for details.

Reminder: the local NEXRADs should be running VCP12.
Next update: usual time, Sunday.

- Martner

December 16, 2005

Project Status: 16 December 2005 (Friday noonish)

IOP-2 to commence soon.

After two weeks of very dry weather, we are short on IOPs. Therefore, we will work the approaching storm, even though it doesn't look impressive now.

We will deploy the research radar crews for a modest storm expected to arrive late tomorrow or early Sunday. The x-band radar crew is already on site; SMART-R crew will fly out from Oklahoma tomorrow. At this point we anticipate that radar operations might begin as early as 6 p.m. PST on Saturday, but probably will commence later. The radar crews should plan on staying in California for a few days, because a second storm (bigger but less certain) is in the forecasts for the middle of the week. See "Forecast for details". Soundings from Slough House will not begin until after the holiday break, due to a manpower shortage. Supplemental soundings from Oakland also will not begin until after the break.

Meanwhile, good progress was made during this dry week on instrument setup. Several sites will get comm lines hooked up today and their realtime data will soon be viewable on the HMT06 web page. Storms during the holiday break (22-28 Dec) will not be worked by the deployable crews, but will, of course, be monitored by HMT's numerous unattended instruments.

- Martner

December 15, 2005

Project Status: 15 December 2005 (Thursday afternoon)

Shorter term forecast indicates dry weather will continue into Saturday. Then a fairly weak storm is expected to drop up to 1 inch of precip (probably rain) on Sunday. Beyond that, next Tues-Weds may be the beginning of a very wet period just before our holiday break. See "Forecast" for details.

Sunday's minor storm may be worth a deployment, considering it might be followed 2-3 days later by a big storm. The holiday air travel crush will make this more interesting, of course. The SMART-R radar crew and the ESRL sounding crew should start getting ready for a possible deployment, with travel to California on Saturday (Dec. 17) and perhaps not returning home until Thursday, Dec. 22. I will make the call on the California-bound travel after tomorrow's conference call update. The X-band crew will stay put in Auburn for now.

-Martner

December 14, 2005

Project Status: 14 December 2005 (Wednesday afternoon)

Dry weather is expected to prevail in the ARB through about Saturday. After that, some models begin to take a little more optimistic trend. SMART-R radar and ESRL sounding crews will remain home for the time being, but should keep in mind the possibility of working a storm shortly before the hoilday break. Next update will be tomorrow.

Tests with the X-band radar are in progress this week at Auburn and ESRL engineers are busily setting up the remainder of the HMT instruments.

- Martner

Project Status: 13 December 2005 (Tuesday night)

No IOPs are ongoing. The west coast ridge seems to be getting locked in place. There will not be any IOP's for the next few days.

Jorgensen

December 12, 2005

Project Status: 12 December 2005 (Monday afternoon)

Forecasts are getting more and more pessimistic. Looks like chances for precip in the ARB are nearly zero through Friday of this week. That will be good for the ESRL engineers while they set up more HMT instruments. SMART-R research radar crew will remain in Oklahoma for the time being. Martner will travel to Auburn tomorrow to check on the X-band radar and hunt for a new disdrometer site. Dave Jorgensen will lead tomorrow's conference call.

- Martner

December 11, 2005

Project Status: 11 December 2005 (Sunday afternoon)

Short-term forcasts call for cloudiness and perhaps sprinkles tomorrow in the ARB followed by dry weather Tuesday through Wednedsaday or Thursday. Beyond that, the models disagree. Yesterday's very optimistic long-range GFS predicition is not looking nearly as promising now for storminess at the end of this week. See "Forecast for details. The reasearch radar crews will remain at their home base and wait for clearer indications of storms moving into California. ESRL's engineers are hitting the road today from Colorado to begin the final round of instrument installments for HMT.

- Martner

December 10, 2005

Project Status: 10 December 2005 (Saturday afternoon)

Forecasts call for dry weather in N. Calif. through about the middle of the upcoming week. The GFS's longer range progs, however, have taken a turn toward big storminess for the latter part of the week (see "Forecast" for details). Let's hope that occurs, but skepticism is warranted until the lead time is shorter. For now, the research radar crews can stay home and stay tuned. Next, update will be tomorrow.

- Martner

December 09, 2005

Project Status: 9December 2005 (Friday afternoon)

The operational forecast models all seem to agree that there will be no significant precipitation in the ARB area before the latter part of next week. Such a dry spell would work to the advantage of the ESRL engineers who will be installing the remainder of the HMT instrumentation next week, including the balloon sounding system. Based on the forecast, the Oklahoma and Colorado radar crews can remain home this weekend. Regardless of the weather, I plan to fly out to Auburn on Tuesday, Dec 13, to attend to some things at the X-band radar and to begin searching for a better site for one of the raindrop disdrometers. Conference calls will continue, as usual, this weekend.

- Martner

December 08, 2005

Project Status: 8 December 2005 (Thursday afternoon)

Crews for the research scanning radars are currently at their home bases in Oklahoma and Colorado while dry weather moves into northern California. Yesterday’s minor event brought only 0.3 inches or less to the ARB, which was not significant enough to warrant deployment. Hence, the earlier decision not to deploy this week was a good one.

The operational models show dry weather dominating through this weekend and well into next week. The teleconnections view across the Pacific, however, holds out hope that the operational models are missing activity that could move into California by Tuesday. See “Forecast” link for details. Meanwhile, the crews will sit tight and wait for clearer indications on tomorrow’s conference call.

The HMT-06 web page “High Resolution Models” link is now up to NOAA/ESRL/GSD’s products for HMT.

Today’s conf call participants: NWS: Western Region, Sacramento RFC, Sacramento WFO, Reno WFO, Monterey WFO, HPC, NOAA/NSSL, and NOAA/ESRL.

- Martner

Status Report: 7 December 2005

Sorry for the late status report. Dry weather expected over the ARB during the weekend. Hoping for some more promising weather next week.

Kingsmill

December 06, 2005

Status Report: 6 December 2005

The event for Wednesday and Thursday continues to look somewhat mediocre, with a poor setup for orographic enhancement. As was stated yesterday, we will not deploy crews to sample this event. It appears that our next best chance for a significant precipitation event is not until the middle of next week.

Kingsmill

December 05, 2005

Status Report: 5 December 2005

The forecast continues to call for a low-QPF event in the American River Basin area during the Wednesday-Thursday time frame. However, given the weakness of the event and the low probability of precipitation in the days after the event, we will not deploy crews to our target area to collect data.

A summary for IOP 1 is now on the Operations Summary blog.

Kingsmill

Status Report: 4 December 2005

The forecast calls for the possibility of a low-QPF event in the American River Basin during the Wednesday-Thursday time frame of this week (7-8 December). However, forecast confidence in this scenario is not particularly high. At this point, we do not anticipate activating an IOP for this event. However, we will monitor its evolution to determine if a change in strategy is warranted.

Kingsmill

December 04, 2005

Status Report: 3 December 2005

Crews returned to their home bases today. As the forecast indicates, dry weather appears to be on the horizon in our target area for at least the next several days. We will continue to monitor the situation on a daily basis.

Kingsmill

December 03, 2005

Status Report: 2 December 2005

The X-POL and SMART-R radars operated until early this morning. Based on the outlook for dry weather over the next several days in our study area, the crews on station in the American River Basin will be returning to their home bases Saturday.

December 01, 2005

Status Report: 1 December 2005

Bursts of precipitation continue to affect the American Basin. X-POL and SMART-R continue to operate and collect data. They are attempting to synchonize their scanning as best as possible with the VCP 12 scans of KDAX and KBBX. We anticipate radar operations to continue all thru the day and into tonight, perhaps terminating after local midnight sometime.

Here is an update on the status of some HMT-West 2006 observing systems:

SMART-R: operating and collecting data
X-POL: operating and collecting data---data is being telemetered to Boulder---working on displaying the data on the web.

Blue Canyon (BLU): data logger battery failed last night---system was back up at 22 UTC
Colfax (CFX,): All data systems (sfc met, precip gauge, soil moisture) are recording, the data is being telemetered back to Boulder and is being displayed on the web site
Colfax (CFC): S-band profiler is not functioning--problems need to be addressed back at the lab--will attempt to deploy the repaired system during the week of 12 December. Other data systems (sfc met, precip gauge, disdrometer) are recording. We are awaiting installation of a phone line at the site to telemeter data to Boulder.
Alta (ATA): All data systems (S-band profiler, sfc met, precip gauge, soil moisture) are recording. We are awaiting installation of a phone line at the site to telemeter data to Boulder.
Foresthill (FHL): All data systems (sfc met, precip gauge, soil moisture) are recording. We are awaiting installation of a phone line at the site to telemeter data to Boulder.
Sloughhouse (SHS): 915 Profiler, GPS IWV, GPS Balloon sounding system, sfc met, rain gauge are scheduled to be installed during the week of 12 December
Yankee Jim Crossing (YJC): Sfc met and rain gauge are scheduled to be installed during the week of 12 December
Meadow Vista (MVA): Installation of disdrometer, sfc met and rain gauge will occur after site approval logistics have been completed (hopefully, by the end of December or early January)

Kingsmill